OpenStudy (dls):

If P(-3,2) is one end of the focal chord PQ of the parabola y^2+ 4x + 4y =0, then the slope of the normal at Q is: (A) -1/2. (B) 2. (C) 1/2. (D) -2

OpenStudy (shubhamsrg):

Whats your progress ?

OpenStudy (dls):

\[\LARGE (y+2)^{2}=4(1-x) \] so vertex=(1,-2)

OpenStudy (dls):